OpenStudy (anonymous):

Is this RIGHT ? Which of the following is an example of the inverse of an exponential function? A. y = B. y = 4x C. y = 2x MY CHOICE! D. y = log2x

OpenStudy (anonymous):

Recheck your choice. It might be correct, or it might be incorrect.

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

do you know what an exponential function looks like?

OpenStudy (anonymous):

The inverse must include the log. or ln

OpenStudy (anonymous):

right, so if you want the inverse you need the natural log.

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

And exponential function is \(f(x) = a^x\) for some number a . The inverse of that will be \(log_a{x}\)

OpenStudy (anonymous):

Was that supposed to say \(\Large \log_2{64} = 6\)?

OpenStudy (anonymous):

Correct. \(2^6 = 64\)

OpenStudy (anonymous):

If I answer do i get a medal?

OpenStudy (anonymous):

Its the same thing

OpenStudy (anonymous):

Best response is the same as a medal, but you shouldn't give best response to someone who tells you the answer. You should give it to someone who teaches you how to find the answer on your own.

OpenStudy (anonymous):

To find the inverse of a function, write a new equation with x and y switched, then solve for y.

OpenStudy (anonymous):

The switching isn't quite like that. You moved the 4 over. When I say switching, I mean, wherever there was an x, write y. Wherever there was a y, write x. Don't change anything else.

OpenStudy (anonymous):

It would be y=(x+2)/4. You achieve this be replacing the x value with y and the y value with x. x=4y-2

OpenStudy (anonymous):

Yes, like that =) Now, just solve your new equation for y. Let me see your attempt.

OpenStudy (anonymous):

I'm sorry, I can't make sense of what you're doing there. That looks like a list of possible answers.

OpenStudy (anonymous):

The switching part gives you x = 4y -2 Now solve that for y. Get y by itself. That means you have to move the 2 over, and then move the 4 over.

OpenStudy (anonymous):

I don't abide guessing.

OpenStudy (anonymous):

Poor you :c

OpenStudy (anonymous):

I can tell you that it is incorrect. Remember your algebra and solve for y. Simple stuff.